OpenStudy (anonymous):

I'm having trouble solving this question, can anyone help me? Find the equation of the tangent line at the given point. f(x)=csc(x) (8, csc(8))

OpenStudy (anonymous):

Find the first derivative of your function. Then evaluate at your point.

OpenStudy (anonymous):

I have my derivative = -csc(x) cot(x) Am I on the right track?

OpenStudy (anonymous):

Yes. -cotxcscx

OpenStudy (anonymous):

Then plug in the x=8 and ill get the y-intercept right?

OpenStudy (anonymous):

Use point slope form.

OpenStudy (anonymous):

\[y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

I've got y=-xcsc(x)cot(x)+8csc(x)cot(x)+csc(8) This doesn't seem right

OpenStudy (anonymous):

No, you need to find the slope at your point first. Do not plug any values into the form yet.

OpenStudy (anonymous):

\[f(x)=cscx\]

OpenStudy (anonymous):

\[f'(x)=-cotxcscx\]

OpenStudy (anonymous):

your point is (8, csc(8))...So..

OpenStudy (anonymous):

\[f'(8)= -\cot(8)\csc(8)\]

OpenStudy (anonymous):

Sorry your last four messages say math processing error, can you input it again please

OpenStudy (anonymous):

\[y-y_{1}= (-\cot(8)\csc(8))(x-x_{1})\]

OpenStudy (anonymous):

f(x)=cscx f′(x)=−cotxcscx your point is (8, csc(8))...So.. f′(8)=−cot(8)csc(8) y−y1=(−cot(8)csc(8))(x−x1)

OpenStudy (anonymous):

y-csc(8)=(-cot(8)csc(8))(x-8) y=(-cot(8)csc(8)x)+8cot(8)csc(8)+csc(8)

OpenStudy (anonymous):

Sorry I was working it out, nice to see I'm on the right track though so far

OpenStudy (anonymous):

Np.

OpenStudy (anonymous):

I've got y=-51.12604x+416.19363

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

A solution and plots using the Mathematica 9 Home Edition program is attached.